QUESTION B1

A study of corruption for a certain geographic region showed an average of 5 corruptions occur per 20,000 people. In a city of 80,000 people, find the probability that at least 3 corruptions occur.

QUESTION B2

(a) A random sample of 300 students from a geographic area indicated that 45 students attended private school. Estimate the true proportion of students attending private schools with 95% confidence.

(b) A survey of 25 individuals who passed the ACCA exam showed that average annual
salary is RM150,000 and standard deviation RM15,000.

(i) Find the degrees of freedom.

(ii) Construct 99% confidence interval for the true average salary.
1
Expert's answer
2020-10-26T19:33:19-0400

QUESTION B1

Let "X="the number of corruption occuried : "X\sim Po(\lambda)"


"P(X=x)=\dfrac{e^{-\lambda}\lambda^x}{x!}"

Given

"\lambda=5\cdot\dfrac{80000}{20000}=20"

"P(X\geq3)=1-(P(X=0)+P(X=1)+P(X=2))="

"=1-\dfrac{e^{-20}\cdot 20^0}{0!}-\dfrac{e^{-20}\cdot 20^1}{1!}-\dfrac{e^{-20}\cdot 20^2}{2!}="

"=1-e^{-20}(1+20+200)=1-221\cdot e^{-20}\approx"

"\approx0.99999954"

QUESTION B2

(a)

Favorable Cases "x=45"  

Sample Size "n=300"  

The sample proportion "\hat{p}=\dfrac{x}{n}=\dfrac{45}{300}=0.15"

The critical value for "\alpha=0.05" is "z_c=1.96". The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}, \hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})=""=(0.15-1.96\sqrt{\dfrac{1.5(1-1.5)}{300}}, 0.15+1.96\sqrt{\dfrac{0.15(1-0.15)}{300}})="

"=(0.1096, 0.1904)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.1096<p<0.1904," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.1094, 0.1904)."


b)

(i)

"df=n-1=25-1=24"

(ii) Based on the provided information, the critical t-value for "\alpha=0.01" and "df=24"

degrees of freedom is "t_c=2.79694."

The 99% confidence for the population mean "\mu" is computed using the following expression


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})="

"=(150000-2.79694\times\dfrac{15000}{\sqrt{25}}, 150000+2.79694\times\dfrac{15000}{\sqrt{25}})="

"=(141609,158391)"